A ship moves from the port of Monrovia to the port of greenville and from there to the port of Buchanan. The first displacement vector is 10.2km due east, and the second is 53.9km due north. What is th resultant displacement vector?



1
Expert's answer
2020-10-05T10:55:34-0400

The magnitude:


"s=\sqrt{s_1^2+s_2^2}=\sqrt{10.2^2+53.9^2}=54.9\ km"

The direction:


"\theta=\arctan{\frac{10.2}{53.9}}=10.7\degree \text{east of north}"